# Rethinking Causal Discovery Through the Lens of Exchangeability

### v1

@@ 4.1 Motivation and Objectives
The creation of this synthetic dataset of cause–effect pairs fulfills two objectives simultaneously:
  • Provide an additional, valid dataset for testing new methods: as noted in Subsection 2.2, the Tübingen dataset remains the sole real-world benchmark for causal discovery. Existing synthetic datasets, in contrast, are used only to evaluate method validity under particular assumptions. The goal of our dataset is to offer a properly constructed synthetic resource for method analysis. We propose using it complementarily alongside the Tübingen dataset: since it is synthetic, it allows for far greater precision (examples can be generated as needed) and excludes certain questionable instances (see Subsection 3.2.2 regarding time-series examples and independence considerations).
  • Showcasing the central role of exchangeability in causal discovery: this dataset is created with a focus on exchangeability. This is achieved by generating samples according to the Causal de Finetti theorem (see Subsection 2.3). With the goal of ensuring exchangeability takes central stage in our dataset, we randomize the choice of both prior distributions and causal functions across different examples, using widely accepted classes. Therefore, by doing so, we aim at showing that if this dataset is able to mimic the real-world dynamics, then exchangeable (not i.i.d.) distributions must be at the center of causal discovery.
@@ 4.2 Algorithm
The algorithm behind the synthetic dataset generation is based on the Causal de Finetti theorem (see Subsection 2.3). According to Equation 8, in order to construct a bivariate dataset satisfying exchangeability, one needs to model the distribution of both latent variables ($p(\theta),p(\psi)$) and the causal functions which are represented by the two conditional probabilities ($p(y_{n}\mid x_{n},\psi),p(x_{n}\mid\theta)$). Given that $x_{n}$ only depends on $\theta$, and the generative nature of our objective, the conditional function relating the two can be fully expressed by taking $\theta$ to represent the true distribution of x, and the conditioning only adding noise to the measurement. On the other hand, $p(y_{n}\mid x_{n},\psi)$ is decomposed into the true causal function $y_{true}=f(x_{n},\psi)$ and a Gaussian noisy measurement $y_{n}=\mathcal{N}(y_{true},\sigma^{2})$ [^9] Note that in both conditional probabilities, we can force the noise power to be zero ($\sigma^{2}=0$) which would yield a noiseless version of the same exchangeable system., with variance $\sigma^{2}$. Therefore, a causal exchangeable dataset generation process can be implemented using Algorithm 1.
@@
Input: $p(\theta),p(\psi),f(),N_{S},\sigma_{x},\sigma_{y}$
   Output: $N_{S}$ causally related exchangeable pairs $\{(x_{i},y_{i})\}$
[…]
2:Perform scaling on both $\theta$ and $\psi$.
@@
[…]
6:Perform min-max scaling on both $X$ and $Y$.
Additionally, to generate many examples, it suffices to wrap Algorithm 1 within an outer loop. In the current implementation, the following considerations are made over the general algorithm and the required hyperparameters and functions (further details about specific design choices made for each causal function can be found in Appendix B.1.)
  • The priors $p(\theta)$ and $p(\psi)$ are chosen at random from 3 different distributions: Uniform, Normal, and Rayleigh. These distributions were chosen because: 1) each distribution captures common patterns observed in empirical data (e.g., a roughly uniform spread of ages in a population or the annual temperature profile approximating a Gaussian). 2) Uniform, Normal, and Rayleigh distributions form foundational building blocks across fields—from entropy-based analyses in information theory to error modeling in signal processing—and underpin key results such as the Central Limit Theorem. Their ubiquity ensures that conclusions drawn under these priors can be readily compared and contextualized with existing work. 3) Once each distribution is linearly scaled to the interval $[0,1]$, no additional shape or scale parameters remain to be tuned, thereby simplifying posterior sampling without introducing extra hyperparameter dependencies.
  • The function $f$ should model common causal relationships in the real-world. Therefore, so far, eight different functions are implemented: 1) linear; 2) piecewise linear; 3) exponential; 4) logarithm; 5) inversely proportional; 6) Brownian-like motion  [^10] Note that this function, despite requiring knowledge of previous samples to compute the latter, does not constitute a timeseries, given there exists no underlying time variable, modeling both $x$ and $y$. Alternatively, it only aims at generating a differentiable function in its most general form.; 7) polynomial; 8) power law. It is also important to point out that all these functions are designed to be strictly increasing in the latent variable ($\psi$). Although this is not directly enforced from the theory of exchangeability, it constitutes a design choice based on theoretical and practical reasons. On the one hand, monotonic SCMs have been proposed as a reasonable simplification of the problem of causal discovery (Izadi and Ester, 2024). On the other hand, many real-world examples in the Tübingen collection (see Appendix A) reinforce this belief. Examples of this are the relationship between age and height, or monthly rent and size in square meters [^11] In the first example, the latent variable can be considered to be the genes; in this case, it is possible that someone with ”taller” genes is both taller than other people at age 10, 40, 70, and so on. In the second example, considering a latent variable such as location, it is also reasonable to assume that independently of the size in square meters of an apartment, the better its location, the more expensive the price..
  • The variances $\sigma_{x}^{2}$ and $\sigma_{y}^{2}$ are small because they are designed to represent noise present in the system.
  • The number of points in each example is sampled from a distribution that aims to mimic the distribution present across the Tübingen dataset. This is obtained by fitting a Gaussian mixture with 3 components to the data.
@@ 4.3 Constructing the Full Dataset
When constructing the full dataset, there exists an additional aspect to consider. In Algorithm 1, three components have been left as design options: the choice of function $f$ and the a priori distributions of the latent variables $p(\theta)$ and $p(\psi)$. Although any fixed choice of these elements yields an exchangeable synthetic dataset, the appropriate frequency of each specification within the global synthetic dataset should be adjusted to mirror the global structure of real data.
To address this issue, nine different established causal discovery methods were implemented: ANM (Peters et al., 2010), bQCD (Tagasovska et al., 2020), CGNN (Goudet et al., 2018), EMD (Chen et al., 2014), IGCI (Mian et al., 2023), LiNGAM (Shimizu et al., 2006), PNL (Zhang and Hyvärinen, 2010), RECI (Blöbaum et al., 2018), and SLOPE (Marx and Vreeken, 2017) [^12] The implementations of RECI and IGCI were taken verbatim from Kalainathan and Goudet (Kalainathan et al., 2020). CGNN and ANM are based on the same repository, with minor adjustments to their hyperparameters (for ANM, we replaced its original independence test with the Hoeffding’s D test to improve computational efficiency). The bQCD and EMD methods were sourced from Tagasovska et al., and SLOPE was implemented as described in its original publication (Tagasovska et al., 2020; Marx and Vreeken, 2017). Finally, the LiNGAM and PNL algorithms were obtained from the GitHub repository at https://github.com/ssamot/causality/tree/master.. Afterwards, their performance in two different metrics, AUROC (area under the ROC curve) and accuracy [^13] These two metrics are widely present in the literature, hence their use in this work., was recorded for each combination of the three existing design choices. In Table 1, partial results obtained by these nine methods are reported. In the absence of any prior bias, the generated dataset was uniformly sampled from across the different combinations of design choices mentioned previously. Consequently, the results displayed in the table can be seen as the average obtained for each function across the different combinations of the choice for the latent variables’ distribution.
@@
Table 1: Marginalized AUROC (accuracy) achieved by different state-of-the-art methods in a uniformly sampled exchangeable synthetic dataset.
Since the algorithms should perform similarly in this dataset as with the Tübingen dataset, in order for it to fulfill its objectives (see Subsection 4.1), the frequency of each combination of parameter choices should be determined such that the results obtained in the two different datasets are as similar as possible. Alternatively, instead of attributing different frequencies to each combination, each one was attributed a different weight (assuming a uniformly sampled distribution across all combinations, i.e., all combinations have the same number of examples). This provides extra flexibility, since different weights can be chosen with different objectives in mind, without the need to generate different examples. Specifically, two different weight combinations are provided in our implementation: one aimed at minimizing the difference in performance with respect to AUROC, the other with respect to accuracy. Both were computed by framing the problem using least squares optimization. In other words, it consists of finding the weight combination that minimizes the sum of the squared difference between each method’s performance in both datasets. Additionally, an $\ell_{2}$ regularizer was also used in order to improve performance for unseen methods [^14] The $\ell_{2}$ regularizer and its weight ($1$) were chosen to optimize the results of SynthNN (see Section 5). In other words, the performance of SynthNN was used as a proxy for the representativeness of the synthetic dataset.. This can be translated into the following mathematical expression:
@@
where $A\in\mathbb{R}^{m\times d}$ is a matrix where element $A_{i,j}$ corresponds to the performance of method $i$ in the generated dataset according to the design choices $j$; $b\in\mathbb{R}^{m}$ is a vector containing the performance of the assessed methods in the Tübingen dataset; $w\in\mathbb{R}^{d}$ is a vector containing the corresponding weight of each parameter combination.
Lastly, AUROC is a ranked choice metric, which means it analyses the order of confidence between different guesses. Therefore, the equation above does not exactly optimize the average error between datasets in the AUROC metric. This is because the weighted average of the AUROC for each design choice combination isn’t the AUROC of the weighted dataset. Nevertheless, optimizing the AUROC exactly is a combinatorial problem, much harder to solve, making the expression above a useful and simple approximation.
@@ 4.4 Analysis
In order for the two objectives laid out initially to be fulfilled, the algorithms should perform similarly in the synthetically generated dataset presented in this paper as in the Tübingen dataset.
Firstly, it is important to understand which shapes the different generated examples can take. The different causal functions presented in Subsection 4.2 were designed to be as general and representative of the real world as possible. Figure 2 shows how synthetic examples can match those in the Tübingen dataset by fixing its random operations. Therefore, it seems reasonable to conclude that the data‑generation mechanism is sufficiently expressive to capture the variety of relationships found in the Tübingen dataset. Nevertheless, despite the ability of the algorithm to generate examples mimicking those in the Tübingen dataset, the inherent randomness present in the process ensures the actual examples are, in practice, much different (a random selection of these examples can be seen in Appendix B.2). It is also important to note that despite this apparent ability to represent the variety of causal functions present in the real-world, the Tübingen dataset has examples with discrete, categorical, or multi-dimensional variables, which are not present in the current implementation of the synthetic dataset.
Figure 2: Three normalized Tübingen pairs plotted alongside hyperparameter-tuned samples from the synthetic dataset proposed in this paper.
Even though the proposed data generation process seems to have the ability (through fine-tuning) to capture real-world examples, this is not enough to guarantee the quality of the dataset for two different reasons. For one, even though the generation process clearly passes the human visual test in terms of expressibility, this does not ensure it provides a proper representation of the causal element inherent to the real-world data. Secondly, even if the algorithm was demonstrably general enough to represent the causal dynamics of all possible known real-world pairs via fine-tuning, this would not necessarily extrapolate accordingly to randomly generated examples. In other words, the noticed causal element present in the fine-tuned examples might be a consequence of the fine-tuning itself, rather than the merits of the data generation algorithm.
@@
Table 2: Performance of different causal discovery methods (in terms of AUROC and accuracy) in the paper’s dataset, the Tübingen, and five additional synthetic datasets. In the ”Ours” datasets, the extra AUROC in parentheses refers to the weighted average AUROC obtained from Equation 9
In order to compare the results shown in Table 2, the average $\ell_{1}$ and $\ell_{2}$ distances to the Tübingen dataset were computed (which can be seen in Table 3). However, since this metric was used for computing the optimal weight distribution for this paper’s dataset, it would be unfair to just present this metric, given the clear circularity in the argument: 1) The dataset quality can be assessed by how similar different methods perform when compared with the real-world reference (the Tübingen dataset) 2) There are certain design choices in the generation process that can be fine-tuned to ensure that a selection of nine different causal discovery methods have very similar performances in both datasets. 3) Since these nine causal discovery methods have very similar performances in both datasets, the data-generation algorithm (and the consequent synthetic dataset) is shown to be very good. By finetuning the dataset using the performances of some causal discovery methods, there is clear confounding in then using the same methods to assess the general similarity in performance of all causal discovery methods in the two datasets. Therefore, to control for this circularity, the average leave-one-out cross-validated $\ell_{1}$ and $\ell_{2}$ distances for this paper’s dataset were also computed. Essentially, it boils down to (one by one) first computing the weights by considering all but one method, and then analyzing the difference in performance only on the left-out method.
@@
Finally, as can be seen in Table 3, the presented dataset performs consistently better than the four current synthetic datasets present in the literature (CE-Gauss, CE-Net, CE-Cha and CE-Multi). The cross validated original version of the synthetic dataset surpasses all other datasets in the $\ell_{1}$ metrics (AUROC: $0.0775$ and accuracy: $0.0516$) and is only nearly outperformed by the CE-Cha dataset in the $\ell_{2}$ metrics (AUROC: $0.0120$ and accuracy: $0.0062$). However, CE-Cha was built for the cause-effect pairs challenge and contains both artificial and real-world data from the Tübingen dataset, which causes clear confounding in this analysis. Furthermore, in absolute terms, the average difference in performance between our dataset and the Tübingen is also quite small. Therefore, it is clear that the presented synthetic dataset correctly mimics the Tübingen one, and consequently, the real-world dynamics accurately. It is also clear that it does so to a higher degree than all other known synthetic datasets.
Additionally, the ”Noisy” dataset was constructed by applying significant additive and multiplicative noise to the original version, and it yields results that are marginally inferior to those of its unaltered counterpart. This suggests that there is nothing inherently causal in the presence of noise that cannot be captured appropriately by the underlying exchangeability in the dataset. This is even more interesting given the absence of any additive or multiplicative relationship between the cause and the latent variable for all implemented causal functions (see Appendix B.1).
@@
  • By allowing exchangeability to drive its development, we have produced the strongest synthetic dataset to date. This success lends weight to the central thesis of our paper: that causal discovery under the i.i.d. assumption should, in fact, be reframed in terms of exchangeability.

### v2

@@ 4.1 Motivation and Objectives
This section describes the creation of a synthetic dataset of cause-effect pairs, generated by following the hierarchical version of the Causal de Finetti representation theorem, expressed in Equation (18). The creation of this synthetic dataset aims to fulfill the two following objectives simultaneously:
  • $\bullet$
    Showcasing the central role of exchangeability in causal discovery, achieved by generating samples according to the hierarchical version of the Causal de Finetti theorem in Equation (18). By doing so, if this dataset is able to match real-world dependency structures, it serves as further confirmation that exchangeable (and not necessarily i.i.d.) distributions are a core concept for causal discovery.
  • $\bullet$
    Providing an additional, valid dataset for testing new causal discovery methods. As noted in Subsection 2.1, the Tübingen dataset remains the sole real-world benchmark for bivariate causal discovery. Existing synthetic datasets, in contrast, are used only to evaluate method validity under particular assumptions. The goal of our dataset is to offer a properly constructed synthetic resource for method analysis. We propose using it complementarily alongside the Tübingen dataset, with its synthetic nature allowing for far greater precision (since arbitrary numbers of examples can be generated as needed).
@@ 4.2 Algorithm
The algorithm for synthetic dataset generation is based on the proposed hierarchical exchangeable model in Equation (18). Firstly, to ensure consistency of orders of magnitude between variables, $X_{n},Y_{n},\eta_{n},$ and $\gamma_{n}$ were always subjected to min-max scaling after sampling. This aims to minimize a method’s ability to artificially explore asymmetries between the two causal directions, arising from designer biases in the ranges of values that the different variables take.
In accordance with Equation (18), and the subsequent interpretation, $\theta$ and $\psi$ will represent both the shape and the parameters of the unknown distributions of the inner latent variables $\eta_{n},\gamma_{n}$, and the choice of functional mechanisms (how each $X_{n}$ depends on $\eta_{n}$, and how each $Y_{n}$ depends on $X_{n}$ and $\gamma_{n}$, respectively). In this generation algorithm, these two possible contributions of $\theta$ and $\psi$ are decoupled. Therefore, Equations (19) and (20) are rewritten by considering $\theta$ and $\psi$ to be two vectors $\theta=[\theta_{1},\theta_{2}];\,\psi=[\psi_{1},\psi_{2}]$, where each subvector controls one of the two possible impacts of $\theta$ and $\psi$, respectively:
$\int_{H}p(X_{n}|\eta_{n},\theta_{2})\,p(\eta_{n}|\theta_{1})\,d\eta_{n}=p(X_{n}|\theta),\,$ (21)
$\int_{\Gamma}p(Y_{n}|X_{n},\gamma_{n},\psi_{2})\,p(\gamma_{n}|\psi_{1})\,d\gamma_{n}=p(Y_{n}|X_{n},\psi).$ (22)
Furthermore, the hierarchical model in Equation (18) does not place any constraints on the values $\theta$ and $\psi$ can take. Nevertheless, in our algorithm, since $\theta_{1}$ and $\psi_{1}$ will represent the choice of the inner latent variables distribution, they were sampled from a uniform discrete distribution over the available distributions. This choice is an algorithm’s hyperparameter, and we opted for the following three possible distributions: Uniform, Gaussian, and Rayleigh. These distributions were chosen given their ubiquitous presence in different applications of statistics. Additionally, these distributions have the significant advantage that their parameters are invariant to min-max scaling. Therefore, there is no need to select the distribution’s parameters.
Variables $\theta_{2}$ and $\psi_{2}$ control the conditional probabilities of $X_{n}$ and $Y_{n}$. In the case of $\theta_{2}$, all information regarding $X_{n}$ can be directly controlled for by the shape of the distribution of $\eta_{n}$ (enforced by $\theta_{1}$). Therefore, this term is redundant and will not be taken into consideration in the presented algorithm. In the case of $\psi_{2}$, it controls the causal mechanism from $X_{n},\gamma_{n}$ to $Y_{n}$. It simultaneously represents the sampled mathematical family of functional mechanisms and the specific function’s parameters (the combination of these two will be denoted $f_{\psi_{2}}$).
Regarding the mathematical family of functions representing the causal relationship, it is uniformly sampled from the following 1) linear; 2) piecewise linear; 3) exponential; 4) logarithm; 5) inversely proportional; 6) Brownian-like motion [^4] Note that this function, despite requiring knowledge of previous samples to compute the latter, does not constitute a time series, given there is no underlying time variable. In fact, it only aims at generating a differentiable function in its most general form.; 7) polynomial; 8) power law. Additionally, regarding the function’s parameters, these are sampled in accordance with fixed hyperparameters and sampling distributions that depend on the specific function. For further details, see Appendix B.1.
Lastly, two additional terms are included to represent the measurement noise of $X_{n}$ and $Y_{n}$. These are two Gaussian distributions with very small variances ($\sigma^{2}_{X}$ and $\sigma^{2}_{Y}$), to ensure they have little impact on the generation process. Thus, the conditional probabilities $p(X_{n}|\eta_{n},\theta_{2})$ and $p(Y_{n}|X_{n},\gamma_{n},\psi_{2})$ take the form
$p(X_{n}|\eta_{n},\theta_{2})=\mathcal{N}(\eta_{n},\sigma^{2}_{X}),$ (23)
$p(Y_{n}|X_{n},\gamma_{n},\psi_{2})=\mathcal{N}\left(f_{\psi_{2}}(X_{n},\gamma_{n}),\sigma^{2}_{Y}\right).$ (24)
To generate a collection of pairs $(X_{1},Y_{1}),...,(X_{N},Y_{N})$, $\theta$ and $\psi$ are sampled once, at the beginning of the generation process, in accordance with the models in Figure 2. Then, the $N$ samples of $\eta_{n},\gamma_{n},X_{n}$, and $Y_{n}$ are obtained as described in Algorithm 1. The value $N$ is sampled from a distribution that aims to mimic the sample size distribution in the Tübingen dataset. This is obtained by fitting a Gaussian mixture with 3 components to the data.
@@
Input: $\{P_{\theta_{1}}\},\{P_{\psi_{1}}\},\{f_{\psi_{2}}\},N,\sigma^{2}_{X},\sigma^{2}_{Y}$
   Output: $N$ causally related exchangeable pairs $\{(x_{i},y_{i})\}$
1:Choose uniformly from $\{P_{\theta_{1}}\},\{P_{\psi_{1}}\},\{f_{\psi_{2}}\}$.
@@
4:Perform min-max scaling on both $(\eta_{1},...,\eta_{N})$ and $(\gamma_{1},...,\gamma_{N})$
5:For $n=1$ to $N,\;X_{n}\sim\mathcal{N}(\eta_{n},\sigma^{2}_{X})$
6:For $n=1$ to $N,\;Y_{n}\sim\mathcal{N}(f_{\psi_{2}}(X_{n},\gamma_{n}),\sigma^{2}_{Y})$
7:Perform min-max scaling on both $(X_{1},...,X_{N})$ and $(Y_{1},...,Y_{N})$.
All considered mathematical descriptions of the causal mechanism (controlled by $\psi_{2}$) are designed to be strictly increasing in the latent variable ($\gamma_{n}$). Although this is not directly enforced by the theory of exchangeability, it constitutes a design choice based on theoretical and practical reasons. On the one hand, monotonic SCMs have been proposed as a reasonable simplification of the problem of causal discovery (Izadi and Ester, 2024); on the other hand, many real-world examples in the Tübingen collection (see Appendix A) reinforce this belief, such as the relationship between age and height, or monthly rent and size in square meters [^5] In the first example, the latent variable can be considered to be the genes; in this case, someone with ”taller” genes may be both taller than other people at age 10, 40, 70, and so on. In the second example, considering a latent variable such as location, it is also reasonable to assume that independently of the size in square meters of an apartment, the better its location, the more expensive the price..
@@ 4.3 Constructing the Full Dataset
Each run of the algorithm described in the previous subsection generates a syntetic cause-effect example. Each example is composed of N samples, and obtained according to a single realization of the outer variables ($\theta$ and $\psi$). In this subsection, we will design a weighting scheme where the examples are weighted differently, based on similarity criteria with certain desired statistical properties. In our case, we will be aiming to mimic the statistics of the Tübingen dataset, and thus of real-world causal systems. To do so, they must depend on the realized value of the outer variables in each example. Specifically, we will be looking at the shape of the distribution of the inner latent variables (defined by $\theta_{1}$ and $\psi_{1}$) and the chosen family of functions to represent the causal mechanisms (defined by $\psi_{2}$). For simplicity, we will not be taking into consideration the specific function parameters.
Alternatively, this process can also be seen as an attempt to model the prior distributions of $\theta_{1}$, $\psi_{1}$, and $\psi_{2}$ to better match observed patterns. Nevertheless, weighting the different combinations of outer variables allows for more flexibility, since different sets of weights can be used based on both differences in applications and differences in similarity criteria.
To help in estimating the distributions of the exchangeable variables, nine different established causal discovery methods were implemented [^6] The implementations of RECI and IGCI were taken verbatim from Kalainathan and Goudet (Kalainathan et al., 2020). CGNN and ANM are based on the same repository, with minor adjustments to their hyperparameters (for ANM, we replaced its original independence test with the Hoeffding’s D test to improve computational efficiency). The bQCD and EMD methods were sourced from Tagasovska et al., and SLOPE was implemented as described in its original publication (Tagasovska et al., 2020; Marx and Vreeken, 2017). Finally, the LiNGAM and PNL algorithms were obtained from the GitHub repository at https://github.com/ssamot/causality/tree/master.: ANM (Peters et al., 2010), bQCD (Tagasovska et al., 2020), CGNN (Goudet et al., 2018), EMD (Chen et al., 2014), IGCI (Mian et al., 2023), LiNGAM (Shimizu et al., 2006), PNL (Zhang and Hyvärinen, 2010), RECI (Blöbaum et al., 2018), and SLOPE (Marx and Vreeken, 2017). Their performance in two different metrics, AUROC (area under the ROC curve) and accuracy (the two metrics are most widely present in the literature), was recorded for each combination of the three relevant exchangeable variables. In Table 1, marginalized results over uniform choices of $\theta_{1},\psi_{1}$ can be seen for these nine methods.
@@
Table 1: Marginalized AUROC (accuracy) over uniform choices of $\theta_{1},\psi_{1}$ achieved by different state-of-the-art methods.
Given the objective of representing real-world causal systems, the weights of the distributions of the exchangeable variables ($\theta_{1},\psi_{1}$ and $\psi_{2}$) were chosen so that the difference in performance for each method between the generated synthetic dataset and that of the Tübingen was minimized. Moreover, two different weight combinations are provided in our implementation: one aimed at minimizing the difference in performance with respect to AUROC, the other with respect to accuracy. Both were computed by framing the problem using least squares optimization. In other words, it consists of finding the weight combination that minimizes the sum of the squared difference between each method’s performance in both datasets. Additionally, an $\ell_{2}$ regularizer was used to improve performance for unseen methods  [^7] The $\ell_{2}$ regularizer and its weight ($1$) were chosen to optimize the results of SynthNN (see Section 5). In other words, the performance of SynthNN was used as a proxy for the representativeness of the synthetic dataset.. This can be translated into the following mathematical expression:
@@
where $A\in\mathbb{R}^{m\times d}$ is a matrix where element $A_{i,j}$ corresponds to the performance of method $i$ in the generated dataset according to the combination of exchangeable variables $j$; $b\in\mathbb{R}^{m}$ is a vector containing the performance of the assessed methods in the Tübingen dataset; and $w\in\mathbb{R}^{d}$ is a vector representing the minimization variable: the corresponding weights of each exchangeable variables combination.
Lastly, AUROC is a ranked choice metric, which means it analyses the order of confidence between different guesses. Therefore, the equation above does not exactly optimize the average error between datasets in the AUROC metric. The weighted average of the AUROC for each combination of exchangeable variables is not the AUROC of the weighted dataset. Nevertheless, optimizing the AUROC exactly is a combinatorial problem, much harder to solve, making the expression above a useful and simple approximation.
@@ 4.4 Analysis
In order for the two objectives laid out initially to be fulfilled, the causal discovery algorithms should perform similarly in the synthetic dataset presented in this paper and in the Tübingen dataset.
Firstly, it is important to understand which shapes the different generated examples can take. The different causal functions were designed to be as general and representative of the real world as possible. Figure 3 shows how synthetic examples can match those in the Tübingen dataset by fixing the choice of function parameters. It seems reasonable to conclude that the data‑generation mechanism is sufficiently expressive to capture the variety of relationships found in the Tübingen dataset. Nevertheless, despite the ability of the algorithm to generate examples mimicking those in the Tübingen dataset, the inherent randomness present in the process ensures the actual examples are, in practice, more different (a random selection of these examples can be seen in Appendix B.2). It is also important to note that despite the apparent ability to represent a variety of causal functions present in the real-world, the Tübingen dataset has examples with discrete, categorical, or multi-dimensional variables, which are not present in our synthetic dataset.
Figure 3: Three normalized Tübingen pairs plotted alongside hyperparameter-tuned examples of the synthetic dataset proposed in this paper.
Even though the proposed data generation process has the ability (through fine-tuning) to capture real-world examples, this is not enough to guarantee the quality of the dataset. For one, even though the generation process clearly passes the human visual test in terms of expressibility, this does not ensure it provides a proper representation of the causal element inherent to the real-world data. Secondly, even if the algorithm was demonstrably general enough to represent the causal structure of known real-world pairs via fine-tuning, this would not necessarily extrapolate to randomly generated examples. In other words, the clear causal element present in the fine-tuned examples might be a consequence of the fine-tuning itself, rather than the merits of the data generation algorithm.
@@
Table 2: Performance of different causal discovery methods (in terms of AUROC and accuracy) in the proposed syjthetic dataset, the Tübingen dataset, and five additional synthetic datasets. In the ”Ours” datasets, the extra AUROC in parentheses refers to the weighted average AUROC obtained from Equation (25)
In order to compare the results shown in Table 2, the average $\ell_{1}$ and $\ell_{2}$ distances to the Tübingen dataset were computed (which can be seen in Table 3). However, since this metric was used for computing the optimal weight distribution for this paper’s dataset, it would be unfair to just present this metric, given the clear circularity in the argument: 1) The dataset quality can be assessed by how similar different methods perform when compared with the real-world reference (the Tübingen dataset) 2) The distribution of the exchangeable variables in the generation process was tuned to ensure that a selection of nine different causal discovery methods have very similar performances in both datasets. 3) Since these nine causal discovery methods have very similar performances in both datasets, the data-generation algorithm (and the consequent synthetic dataset) is shown to be very good. By finetuning the dataset using the performances of causal discovery methods, there is clear confounding in then using the same methods to assess the general similarity in performance for all causal discovery methods in the two datasets. Therefore, to control for this circularity, the average leave-one-out cross-validated $\ell_{1}$ and $\ell_{2}$ distances for this paper’s datasets were also computed. It boils down to (one by one) first computing the weights by considering all but one method, and then analyzing the difference in performance only on the left-out method.
@@
Finally, as can be seen in Table 3, the presented dataset performs consistently better than four current synthetic datasets present in the literature (CE-Gauss, CE-Net, CE-Cha, and CE-Multi). The cross-validated original version of the synthetic dataset surpasses all other datasets in the $\ell_{1}$ metrics (AUROC: $0.0775$ and accuracy: $0.0516$) and is only nearly outperformed by the CE-Cha dataset in the $\ell_{2}$ metrics (AUROC: $0.0120$ and accuracy: $0.0062$). However, CE-Cha was built for the cause-effect pairs challenge and contains both artificial and real-world data from the Tübingen dataset, which causes clear confounding in this analysis. Furthermore, in absolute terms, the average difference in performance between our dataset and the Tübingen is also quite small. In all, it is clear that the presented synthetic dataset correctly mimics the Tübingen and consequently, real-world dynamics accurately. It is also clear that it does so to a higher degree than all other known synthetic datasets.
Additionally, the ”noisy” dataset was constructed by applying significant additive and multiplicative noise to the original version, and it yields results that are marginally inferior to those of its unaltered counterpart. This suggests that there is nothing inherently causal in the presence of noise that cannot be captured appropriately by the underlying exchangeability in the dataset. This is even more interesting given the absence of any additive or multiplicative relationship between the cause and the inner latent variable ($\gamma_{n}$) for all implemented causal functions (see Appendix B.1).
@@
  • $\bullet$
    By allowing exchangeability to drive its development, we have produced a very strong synthetic dataset to date. This success lends weight to the central thesis of our paper: that causal discovery under the i.i.d. assumption should, in fact, be reframed in terms of exchangeability.